a from a balloon reaches the ground in 30 sec. Determine the height of the balloon of it is ascending with a speed of 100cm/s when the bomb is dropped


1
Expert's answer
2021-01-13T11:37:34-0500

The equation of free fall is:


"h = -v_0t +\dfrac{gt^2}{2}"

where "v_0 = 100cm\/s = 1m\/s" (- sign is due to opposite directions of "g" and "v_0") is the initial speed of the bomp, "t = 30s" is the time of falling, and "g = 9.81m\/s^2" is the gravitational acceleration. "h" is the height of the balloon. Thus, obtain:


"h = -1\cdot 30 + \dfrac{9.81\cdot 30^2}{2} = 4384.5m"

Answer. 4384.5 m.